A right triangle has one leg four times as long as the other. Find a function that models its perimeter P in terms of the length x of the shorter leg.
Answer by stanbon(75887)

A right triangle has one leg four times as long as the other. Find a function that models its perimeter P in terms of the length x of the shorter leg.
--------------------------------------
Let shortest leg be "x"
2nd leg = 4x
hypotenuse = sqrt(x^2 + (4x)^2) = sqrt(17x^2) = x*sqrt(17)
--------------------------------------------------------------
Perimeter = x + 4x + x*sqrt(17)
P = (5+sqrt(17)x
====================================
Cheers,
Stan H.